what is the value of x in the equation 13x-2(x+4)=8x+1

Found 2 solutions by checkley71, oscargut:
Answer by checkley71(8403)

13x-2(x+4)=8x+1
13x-2x-8=8x+1
13x-2x-8x=1+8
3x=9
x=9/3
x=3 answer.
proof
13*3-2(3+4)=8*3+1
39-2*7=24+1
39-14=25
25=25

Answer by oscargut(686)

13x-2(x+4)=8x+1 then 11x-8=8x+1 then 3x=9
then x=3